# CAPability: A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

### v3

@@ 1 Introduction
Visual captioning, which translates visual content into textual descriptions, is a fundamental task for both image and video understanding fang2015captions ; xu2015show , and forms a significant basis for image and video generation peebles2023scalable ; wang2023modelscope . To assess the capabilities of this task, researchers established several visual caption benchmarks in earlier years mscoco ; nocaps ; msrvtt ; vatex .
@@
However, the evaluation of these benchmarks remains incomplete as they focus on a single aspect of captions with limited visual elements, inadequately covering the full caption scope. For instance, they often overlook aspects like scene, text, and style. We argue that a multi-view evaluation for visual caption is essential. In this paper, we introduce a new comprehensive visual caption benchmark, CAPability. Our approach uses complete visual elements rather than caption sentences as annotations to evaluate both correctness and thoroughness for each dimension. The selection of dimensions is motivated by existing visual generation benchmarks geneval ; vbench ; t2vcompbench , which is the inverse task of visual captioning. They usually treat and evaluate the generated visual content from different aspects (e.g., objects, scene, style, camera, motion control). Similarly, we design 6 views and 12 dimensions for CAPability by analyzing several visual captions, as illustrated in Tab. 1 and Fig. 1. We believe these components contribute to a complete caption, as lacking any of them may align the caption with different visual content. There are 9 static and 4 dynamic dimensions, with object number encompassing both static and dynamic aspects. Dynamic dimensions apply to both images and videos, while static ones are exclusive to video. We collect and manually annotate 11K images and videos for CAPability, providing sufficient samples.
@@
Benchmark Data Type Anno- tations Evaluation
@@
In addition to multi-view annotation, we also focus on improving the evaluation quality of captions. While most methods assess only the correctness, we argue that considering both correctness and thoroughness of visual elements provides a more comprehensive evaluation for visual captioning. Therefore, we conduct comprehensive experiments using both precision and hit as our main metric, supplemented by another heuristic metric: the know but cannot tell ($K\bar{T}$). While precision only evaluates the accuracy of captions, hit considers both accuracy and coverage of captions compared with ground truth, and $K\bar{T}$ indicates the ratio when a model can answer related questions correctly but fails to convey the same information in the caption automatically. These metrics provide a robust framework for evaluating both correctness and thoroughness in MLLMs’ detailed captions. To our knowledge, we are the first to heuristically highlight the gap in correctness and thoroughness capabilities of MLLMs across multiple views, guiding researchers to enhance these capabilities across dimensions. Representative results are shown in Fig. 3, leading to the following conclusions: 1) Models perform more variably on hit, which evaluates the thoroughness and may be ignored by previous research. 2) GPT-4o seems the best on precision. For hit, Gemini-1.5-pro gemini1.5 has a leading advantage in many dimensions, followed by GPT-4o gpt4o . 3) Gemini-1.5-pro demonstrates strong object-counting abilities, while GPT-4o excels in identifying camera angles. 4) All models still struggle with dimensions like object numbers, camera angle, camera movement, character identification, and action. We hope our findings guide researchers to focus on improving these abilities in caption tasks. Our main contributions are listed as follows:
@@ 2 Related Work
Multi-modal large language models. Based on the significant development of Large Language Models (LLMs) among various linguistic tasks gpt3 ; vicuna ; qwen2.5 ; llama3 , many works try to extend the powerful capabilities into multi-modal understanding. By integrating image content into LLMs, Multi-modal Large Language Models (MLLMs) also gain huge achievements llava ; llava1.5 ; qwenvl ; minigpt4 ; llavanext ; internvl ; gpt4v . Based on the pre-trained weights from image models, recent MLLMs also expand video understanding capabilities llavanextvideo ; llavaov ; llavavideo ; qwen2vl ; internvl2.5 ; nvila ; videollama3 ; qwen2.5vl . With rapid development, MLLMs are powerful enough to describe both the image and video content in detail, which makes the traditional benchmarks with short captions outdated. More and more methods even try to produce re-captioned detailed descriptions by more powerful models rather than existing human-annotated short captions to train their model llavaov ; videollama3 . Therefore, it is urgent to propose a new visual caption benchmark that adapts to modern MLLMs.
@@ 3.1 Multiple Dimension Data Annotation
The pipeline of our whole collection and annotation is shown in Fig. 4. We first design 6 views and split 12 dimensions, then collect nearly 1,000 images and videos for each dimension separately. For the collected data, we conduct pre-annotations by SOTA MLLMs and the following data balancing before the human annotation. After the human-annotation of each dimension, we filter bad cases during the annotation and finally complete the data of CAPability.
@@
Data collection. For convenience and problem simplification, we only collect image data for static dimensions and video data for dynamic dimensions. This is based on the common sense that the video understanding capabilities for MLLMs are usually built upon sufficient image understanding capabilities llavanextvideo ; llavavideo ; videollama3 ; internvl2.5 . Since an image or a video cannot cover all these dimensions of information, we directly collect data for each dimension independently and evaluate each dimension separately. For static dimensions, we mainly collect images from SA-1B sam , COYO-700M coyo700m , Wukong wukong100m , and Wikipaintings wikipaintings , and we also crawl a considerable amount of data from multiple public datasets and websites with CC0 license by ourselves. We also borrow parts of the image data and annotations from CompreCap comprecap for the spatial relation dimension. For dynamic dimensions, we crawl and cut videos for camera movement dimension, borrow videos from Dream-1K dream1k for action and event dimensions, and borrow videos from VSI-Bench vsibench for the dynamic object number dimension. Fig. 5 shows our data sources for each dimension and their proportion.
@@ 4.2 Main Evaluation Results
Precision and hit of closed-source API and 72B models. We report the precision and hit of closed-source and 72B models in Tab. 3. Gemini-2.0-flash and GPT-4o achieve the highest precision (79.3% and 79.2%), which represents their captions are truthful and accurate. When it comes to the hit metric, all results drop significantly as it is harder to cover as many visual elements as possible. Gemini-1.5-pro achieves the best with 3.9% higher than second place, i.e., GPT-4o, which means it is better at identifying more elements correctly. Among 72B models, Qwen2.5VL achieves the best hit, with notable precision. When we focus on each dimension, it is worth noting that these models behave differently in different dimensions. Gemini-1.5-pro has a huge advantage in object counting in both image and video, especially in thoroughness (10% better than the second place for image and 16.3% better than the second place for video in hit metric). GPT-4o excels at recognizing camera angle, as it is 9.6% higher than the second place on both precision and hit. Qwen2.5-VL performs the best hit in the open-source models as it performs well on scene and camera movement. Object category, scene, OCR, and style seem simple for these powerful models, as they all achieve well on the F1-score. When it comes to the dimensions of object number, object color, spatial relation, style, character identification, and events, all of them show relatively high precision but low hit, which means they can describe these elements well when they are confident about them, but might miss some instances and ignore the thoroughness. As for the action and two camera-related dimensions, all models perform unsatisfactorily on both precision and hit. This phenomenon inspires researchers to focus more on these aspects of the model’s capability.
@@ Methods
Table 5: The $\uparrow$ (higher is better) of CAPability-QA and the $K\bar{T}\downarrow$ (lower is better) result.

### v4

@@ 1 Introduction
Visual captioning, which translates visual content into textual descriptions, is a fundamental task for both image and video understanding, affecting various downstream tasks jiang2023efficient; liu2024towards; jiang2025videopure. The caption capability directly reflects the modality alignment ability fang2015captions; xu2015show, forms a significant basis for image and video generation peebles2023scalable; wang2023modelscope; wei2024dreamvideo; wei2025dreamrelation, and provides the precondition of synthesizing a large-scale multi-modal understanding and reasoning dataset llavavideo; arnab2025temporal. To assess the capabilities of this task, researchers established several visual caption benchmarks in earlier years mscoco; nocaps; msrvtt; vatex.
@@
However, the evaluation of these benchmarks remains incomplete as they focus on a single aspect of captions with limited visual elements, inadequately covering the full caption scope. For instance, they often overlook aspects like scene, text, and style. We argue that a multi-view evaluation for visual caption is essential. In this paper, we introduce a new comprehensive visual caption benchmark, CAPability. Our approach uses complete visual elements rather than caption sentences as annotations to evaluate both correctness and thoroughness for each dimension. The selection of dimensions is motivated by existing visual generation benchmarks geneval; vbench; t2vcompbench, which is the inverse task of visual captioning. They usually treat and evaluate the generated visual content from different aspects (e.g., objects, scene, style, camera, motion control). Similarly, we design 6 views and 12 dimensions for CAPability by analyzing several visual captions, as illustrated in Tab. 1 and Fig. 1. We believe these components contribute to a complete caption, as lacking any of them may align the caption with different visual content. There are 9 static and 4 dynamic dimensions, with object number encompassing both static and dynamic aspects. Static dimensions apply to both images and videos, while dynamic ones are exclusive to video. In our CAPability, we collect video data only for dynamic dimensions and image data for static dimensions for simplicity. We then manually annotate 11K images and videos for CAPability, providing sufficient samples.
@@
Benchmark Data Type
Anno-
tations
Evaluation
@@
In addition to multi-view annotation, we also focus on improving the evaluation quality of captions. While most methods assess only the correctness, we argue that considering both correctness and thoroughness of visual elements provides a more comprehensive evaluation for visual captioning. Therefore, we conduct comprehensive experiments using both precision and hit as our main metric, supplemented by another heuristic metric: the know but cannot tell ($K\bar{T}$). While precision only evaluates the accuracy of captions, hit considers both accuracy and coverage of captions compared with ground truth, and $K\bar{T}$ indicates the ratio when a model can answer related questions correctly but fails to convey the same information in the caption automatically. These metrics provide a robust framework for evaluating both correctness and thoroughness in MLLMs’ detailed captions. To our knowledge, we are the first to heuristically highlight the gap in correctness and thoroughness capabilities of MLLMs across multiple views, providing deeper insight into specific capabilities or limitations of a model, thereby offering actionable guidance for further research and development, rather than just an overall score. Representative results are shown in Fig. 3, leading to the following conclusions: 1) Models perform more variably on hit, which evaluates the thoroughness and may be ignored by previous research. 2) GPT-4o seems the best on precision. For hit, Gemini-1.5-pro gemini1.5 has a leading advantage in many dimensions, followed by GPT-4o gpt4o. 3) Gemini-1.5-pro demonstrates strong object-counting abilities, while GPT-4o excels in identifying camera angles. 4) All models still struggle with dimensions like object numbers, camera angle, camera movement, character identification, and action. We hope our findings guide researchers to focus on improving these abilities in caption tasks. Our main contributions are listed as follows:
@@ 2 Related Work
Multi-modal large language models. Based on the significant development of Large Language Models (LLMs) among various linguistic tasks gpt3; vicuna; qwen2.5; llama3, many works try to extend the powerful capabilities into multi-modal understanding liu2025hybrid; bao2024cores; bao2025dynimg. By integrating image content into LLMs, Multi-modal Large Language Models (MLLMs) also gain huge achievements llava; llava1.5; qwenvl; minigpt4; llavanext; internvl; gpt4v. Based on the pre-trained weights from image models, recent MLLMs also expand video understanding capabilities llavanextvideo; llavaov; llavavideo; qwen2vl; internvl2.5; nvila; videollama3; qwen2.5vl. With rapid development, MLLMs are powerful enough to describe both the image and video content in detail, which makes the traditional benchmarks with short captions outdated. More and more methods even try to produce re-captioned detailed descriptions by more powerful models rather than existing human-annotated short captions to train their model llavaov; videollama3. Therefore, it is urgent to propose a new visual caption benchmark that adapts to modern MLLMs.
@@ 3.1 Multiple Dimension Data Annotation
The pipeline of our whole collection and annotation is shown in Fig. 4. We first design 6 views and split 12 dimensions, then collect nearly 1,000 images and videos for each dimension with low overlap among dimensions. For the collected data, we conduct pre-annotations by SOTA MLLMs and the following data balancing before the human annotation. After the human annotation of each dimension, we filter bad cases during the annotation and finally complete the data of CAPability.
@@
Data collection. For convenience and problem simplification, we only collect image data for static dimensions and video data for dynamic dimensions. This is based on the common sense that the video understanding capabilities for MLLMs are usually built upon sufficient image understanding capabilities llavanextvideo; llavavideo; videollama3; internvl2.5. Since an image or a video cannot cover all these dimensions of information, we directly collect data for each dimension independently and evaluate each dimension separately. For static dimensions, we mainly collect images from SA-1B sam, COYO-700M coyo700m, Wukong wukong100m, and Wikipaintings wikipaintings, and we also crawl a considerable amount of data from multiple public datasets and websites with CC0 license by ourselves. We also borrow parts of the image data and annotations from CompreCap comprecap for the spatial relation dimension. For dynamic dimensions, we crawl and cut videos for camera movement dimension, borrow videos from Dream-1K dream1k for action and event dimensions, and borrow videos from VSI-Bench vsibench for the dynamic object number dimension. Fig. 6 shows our data sources for each dimension and their proportion, and Tab. A1 shows our data overlap among dimensions.
@@ 4.2 Main Evaluation Results
Precision and hit of closed-source API and 72B models. We report the precision and hit of closed-source and 72B models in Tab. 3. Gemini-2.0-flash and GPT-4o achieve the highest precision (79.3% and 79.2%), which represents their captions are truthful and accurate. When it comes to the hit metric, all results drop significantly as it is harder to cover as many visual elements as possible. Gemini-1.5-pro achieves the best with 3.9% higher than second place, i.e., GPT-4o, which means it is better at identifying more elements correctly. Among 72B models, Qwen2.5VL achieves the best hit, with notable precision. We note that the high precision but moderate hit for models like GPT-4o suggests they tend to describe only what they are confident about and avoid parts they are less certain of (line 245). This behavior results in high accuracy for described elements, but at the cost of thoroughness and coverage. When we focus on each dimension, it is worth noting that these models behave differently in different dimensions. Gemini-1.5-pro has a huge advantage in object counting in both image and video, especially in thoroughness (10% better than the second place for image and 16.3% better than the second place for video in hit metric). GPT-4o excels at recognizing camera angle, as it is 9.6% higher than the second place on both precision and hit. Qwen2.5-VL performs the best hit in the open-source models as it performs well on scene and camera movement. Object category, scene, OCR, and style seem simple for these powerful models, as they all achieve well on the F1-score. When it comes to the dimensions of object number, object color, spatial relation, style, character identification, and events, all of them show relatively high precision but low hit, which means they can describe these elements well when they are confident about them, but might miss some instances and ignore the thoroughness. As for the action and two camera-related dimensions, all models perform unsatisfactorily on both precision and hit. This phenomenon inspires researchers to focus more on these aspects of the model’s capability. We acknowledge that this finding highlights well-known open challenges for current MLLMs. For example, accurate object counting requires fine-grained discrimination; some recent methods deepeyes; jiang2025vlm leverage segmentation or external tools to assist reasoning. Camera-related tasks demand a strong visual reasoning ability to infer subtle contextual cues. Character identification relies heavily on high-quality training data to build related knowledge. Action recognition is inherently challenging due to the temporal modeling.
@@ Methods
[…]
@@ 5 Conclusion
Acknowledgment
This work is supported by the National Nature Science Foundation of China (62425114, 62121002, U23B2028, 62232006). We thank Alibaba Group for providing human annotation resources, closed-source model APIs support, and GPU computing resources. We thank all outsourcing annotators’ efforts, and acknowledge the support of the GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC, and USTC super-computing center for providing computational resources for this project.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims presented in the abstract and introduction accurately reflect the contributions and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations in the appendix performed by the authors.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: During the inference process, we set the temperature parameter of the MLLMs to 0 to eliminate randomness and ensure the validity of the theoretical results derived. The proposed metrics are discussed with proper reference in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We introduce the detailed steps to obtain experiment results in the paper and explain the implementation details in the appendix. We provide the code and benchmark data to ensure the result can be reproducible. We also provide the inference and evaluation JSON files in the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided open-accessed code and data link in the submission.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Detailed specifics are provided in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the results of repeating the experiment several times in the appendix to verify the evaluation stability.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the computing resource we use in the appendix, though it is not the matter we care about for proposing a benchmark.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper adheres to the NeurIPS Code of Ethics. All experimental procedures, data handling, and reporting practices were conducted with full compliance to ethical standards, ensuring transparency, integrity, and respect for all stakeholders involved
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We highlight some findings based on our evaluation results, which inspire the researchers. We also discuss the potential positive and negative societal impacts in the appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: As mentioned in the paper, we conduct human filtering when annotating every image and video, ensuring there are no unsafe images or videos released.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper provides appropriate credits and references to the creators or original owners of the assets used. We centrally state the licenses and copyright in the appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have made our benchmark data publicly available, and the link and croissant file are provided in the submission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We ask outsourcing annotators to complete the benchmark construction, but it does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We use LLMs for editing typos and formatting, and use LLMs for pre-annotation, followed by human annotation. In our evaluation, we also employ LLM-as-judge evaluation. However, the core motivation, method, and development do not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.